### v1

# VR-GS: A Physical Dynamics-Aware Interactive Gaussian Splatting System in Virtual Reality

Abstract.
* indicates equal contributions.
Affiliations: ${}^{1}$, ${}^{2}$, ${}^{3}$tah, ${}^{4}$, ${}^{5}$tyle3D Research, ${}^{6}$, ${}^{7}$mazon
As consumer Virtual Reality (VR) and Mixed Reality (MR) technologies gain momentum, there’s a growing focus on the development of engagements with 3D virtual content. Unfortunately, traditional techniques for content creation, editing, and interaction within these virtual spaces are fraught with difficulties. They tend to be not only engineering-intensive but also require extensive expertise, which adds to the frustration and inefficiency in virtual object manipulation. Our proposed VR-GS system represents a leap forward in human-centered 3D content interaction, offering a seamless and intuitive user experience. By developing a physical dynamics-aware interactive Gaussian Splatting (GS) in a VR setting, and constructing a highly efficient two-level embedding strategy alongside deformable body simulations, VR-GS ensures real-time execution with highly realistic dynamic responses. The components of our VR system are designed for high efficiency and effectiveness, starting from detailed scene reconstruction and object segmentation, advancing through multi-view image in-painting, and extending to interactive physics-based editing. The system also incorporates real-time deformation embedding and dynamic shadow casting, ensuring a comprehensive and engaging virtual experience. Our project page is available at: https://yingjiang96.github.io/VR-GS/.
copyright: none
Figure 1. Animal Crossing. Utilizing our system, individuals can engage in intuitive and interactive physics-based game-play with deformable virtual animals and realistic environments represented with 3D Gaussian Splatting.
1. Introduction
The burgeoning growth in commercial VR and MR applications is propelling the quest for high-fidelity 3D content within the Computer Graphics (CG) and Computer Vision (CV) communities. Despite their capabilities of being rendered efficiently in VR, traditional 3D/4D content creation, reliant on 3D modeling tools and game engines, is time-intensive and complex, often beyond the reach of non-expert users. This accessibility barrier limits broader user engagement in high-quality content creation.
Recognizing the advantages and limitations of the traditional graphics pipeline, our goal is to find a modern variant. To enhance the visual quality and ease of creation for non-expert users, we move away from traditional 3D models in graphics pipelines and instead, adopt state-of-the-art radiance field techniques for rendering. In this context, Neural Radiance Fields (NeRF) emerge as a natural choice. Despite its versatility, NeRF’s volume rendering falls short in efficiency for interactive rendering on Extended Reality (XR) devices, which demand high frame rates. Additionally, NeRF’s approach to handling deformations—requiring bending of query rays via an inverse deformation map—is slow (Pumarola et al., 2021). Fortunately, 3D Gaussian Splatting (GS) (Kerbl et al., 2023) has been recently introduced as an efficient and explicit alternative to NeRF. This method not only excels in rendering efficiency but also provides an explicit geometric representation that can be directly deformed or edited. The explicit nature of 3D GS can simplify the direct manipulation of geometry by solving a Partial Differentiable Equation (PDE) that governs the motions of Gaussian kernels. Furthermore, GS eliminates the need for high-fidelity meshes, UV maps, and textures, offering the potential for naturally photorealistic appearances. It is worth noting that many studies have already demonstrated the use of 3D GS for 4D dynamics (Park et al., 2021; Pumarola et al., 2021; Yang et al., 2023) and the integration of physics-based simulations for more realistic 4D content generation (Xie et al., 2023; Feng et al., 2023), as well as for creating animatable avatars (Zielonka et al., 2023; Xu et al., 2023b).
In this paper, we introduce VR-GS, a physics-aware, interactive Virtual Reality (VR) system for immersive manipulation of 3D content represented with Gaussian Splatting (GS). To ensure an interactive experience, we utilize eXtended Position-based Dynamics (XPBD) (Macklin et al., 2016), a highly adaptable and unified physical simulator, for real-time deformation simulation. Direct incorporation of a simulator onto GS kernels presents challenges, as the simulation and rendering processes have distinct geometrical representations. To address this, we construct a tetrahedral cage for each segmented GS kernel group and embed these kernel groups into corresponding meshes. The deformed mesh, driven by XPBD, subsequently guides the deformation of the GS kernels. Noticing that simplistic embedding techniques can lead to undesirable, spiky deformations in GS kernels, we propose a novel two-level embedding approach. This method allows each Gaussian kernel to adapt to a smoothed average deformation of the surrounding tetrahedra. The intricate combination of GS and XPBD through our two-level embedding not only achieves real-time physics-based dynamics but also upholds high-quality, realistic rendering, enabling high-fidelity immersive VR experiences. In summary, our contributions include:
  • Immersive VR System Development: The development of a VR system blending physical dynamics with high-fidelity appearance for enriched user experiences. We extensively evaluate our framework with various experiments including user studies.
  • Real-Time 3D Content Interaction: The engineering of the system with a focus on human-centric, creative interaction with captured and virtual 3D content.
  • Comprehensive System Integration: The comprehensive combination of 3D GS, segmentation, image inpainting, and real-time physics-based solvers with a novel rendering geometry embedding algorithm.
2. Related Work
2.1. Radiance Fields Rendering
A variety of 3D representations, such as mesh (Sorkine and Cohen-Or, 2004), point clouds (Qi et al., 2017), signed distance fields (Wang et al., 2021), and grids (Zhu and Bridson, 2005) were exploited in early work for visual computing tasks, including synthesis, estimation, manipulation, animation, reconstruction, and transmission of data about objects and scenes (Kerbl et al., 2023; Xie et al., 2022; Gao et al., 2023). Since neural radiance fields (NeRF) (Mildenhall et al., 2021) was proposed for novel view synthesis with differentiable volume rendering, it has been applied to versatile computer graphics and computer vision tasks, such as scene reconstruction (Oechsle et al., 2021; Meng et al., 2023; Yariv et al., 2021), synthesis (Sun et al., 2023; Huang et al., 2023a), rendering (Lombardi et al., 2021; Wu et al., 2023) and animation (Chen et al., 2021; Peng et al., 2021), simulation (Li et al., 2023a), etc. While Neural Radiance Fields (NeRF) have achieved remarkable image quality, they suffer from significant time and memory inefficiencies (Lindell et al., 2021). To mitigate these issues, various methods have been introduced. Sparse representations (Liu et al., 2020), decomposed strategies (Niemeyer and Geiger, 2021; Rebain et al., 2021), and multi-resolution encoding (Müller et al., 2022) have all been proposed to enhance volume rendering efficiency while maintaining high quality. However, NeRFs are implicit representations, making it challenging to detect and resolve collisions (Qiao et al., 2023). Recently, 3D Gaussian splatting (GS) proposed by Kerbl et al. (2023) utilizes an array of 3D Gaussian kernels to represent scenes explicitly, facilitating faster optimization and achieving state-of-the-art results.
2.2. Editing in Radiance-based Scene
A natural extension of NeRF is to support user-guided editing. Li and Pan (2023) made use of two proxy cages to offer interactive control of the shape deformation of NeRF. Besides geometry editing, text prompts (Wang et al., 2023) or a user-edited image (Bao et al., 2023) can also be adopted to conduct style transfer of NeRF. Besides, Lin et al. (2023) put forward exploiting 2D sketches to edit high-quality facial NeRF. Concerning NeRF texture editing, Huang et al. (2023a) utilized a coarse-fine disentanglement representation and a patch-matching algorithm to synthesize textures of different shapes from multi-view images. What’s more, De-Nerf (Wu et al., 2023) exploits a hybrid light representation that enables users to relight NeRF. In contrast, Gaussian Splatting is more appropriate for post-editing tasks thanks to its explicit nature and inspired a series of follow-up works (Chen et al., 2023; Fang et al., 2023; Ye et al., 2023; Duisterhof et al., 2023; Huang et al., 2023b).
Another significant direction in NeRF research is the incorporation of dynamic details into static scenes. To achieve this, a popular strategy (Pumarola et al., 2021; Park et al., 2021) is to consider time as an additional input to the system, This method typically splits a dynamic NeRF into a canonical static field and a deformation field, with the latter mapping this canonical representation to a deformed one. More recently, Yang et al. (2023) drew inspiration from 3D GS to reconstruct a dynamic scene using 4D Gaussian primitives that change over time. However, the dynamics in these methods are generally confined to motions captured from input data, limiting their ability to synthesize unseen dynamics. To generate novel dynamics, Xie et al. (2023) and Feng et al. (2023) have integrated physics-based simulations into 3D static representations, paving the way for generating physically plausible and novel dynamic scenes.
2.3. Neural Radiance Fields in VR
Rendering NeRF is computationally expensive for real-time VR applications, which causes high latency and low quality (Song et al., 2023; Li et al., 2022b). To address these challenges and enable high-fidelity rendering with minimal latency on a single GPU, various techniques have been proposed. These include gaze-contingent 3D neural representations (Deng et al., 2022), variable rate shading (Rolff et al., 2023), and hybrid surface-volume representations (Turki et al., 2023), all aimed at accelerating NeRF. In contrast, VR-NeRF (Xu et al., 2023a) leverages multiple GPUs to achieve high-quality volumetric rendering of NeRF. Beyond software acceleration techniques, RT-NeRF (Li et al., 2022a) utilizes an algorithm-hardware co-design framework to provide real-time NeRF solutions for immersive VR rendering. While these methods primarily focus on enhancing NeRF rendering, our work is dedicated to facilitating interactive editing within a VR environment.
Li et al. (2023c) integrated affine transformation with NeRF to enable interactive features, such as exocentric manipulation, editing, and VR tunneling effects. However, their approach is limited to filter-based edits such as edge blurring and fails to support the deformation of virtual objects’ geometry. RealityGit (Li et al., 2023b) and Magic NeRF Lens (Li et al., 2023d) modify a NeRF model by erasing or revealing a portion of NeRF. Nevertheless, neither system possesses the functionality to deform the geometrical structure of the model as well. In contrast to these works which concentrate on interactive editing of NeRF, our paper introduces VR-GS, a new physical dynamics-aware interactive system for GS in VR, first enabling users to interact with 3D GS at a real-time rate.
3. Design Goals and Design Decisions
We propose the development of a unified, physics dynamics-aware interactive VR system, meticulously designed for real-time interaction with Gaussian Splatting (GS).
Figure 2. VR-GS is an interactive system meticulously crafted to integrate 3D Gaussian Splatting (GS) and eXtended Position-based Dynamics (XPBD) for generating a real-time interactive experience in Virtual Reality (VR). Beginning with multi-view images, the pipeline skillfully combines scene reconstruction, segmentation, and inpainting using Gaussian kernels. These kernels form the foundation for VR-GS’s utilization of the sparse volumetric data structure VDB, facilitating bounding mesh reconstruction and subsequent tetrahedralization. VR-GS further harnesses a novel two-level rendering geometry embedding, XPBD, collision detection, and shadow casting techniques, all converging to deliver a captivating and immersive user experience.
3.1. Design Goals
G1. Immersive and Realistic
A VR system should provide a deeply immersive 3D environment that convincingly simulates physical presence, where objects possess a high degree of realism (Kalawsky, 1999). Accordingly, our targeted VR system aims to offer users an experience that is both immersive and realistic, characterized by dynamic motions, 3D virtual shapes, and illuminations that closely mirror the real world. This goal sets our system apart, especially in how it handles physics-based 3D deformations and dynamics, as opposed to those that rely solely on geometric deformation.
G2. Real-Time Interaction
For a VR system, low latency is crucial to prevent disorientation or motion sickness (Sutcliffe et al., 2019; Deng et al., 2022). Slow system responses can severely impair the quality of interaction. Our proposed VR system is engineered to provide users with instantaneous output and feedback during their interaction. This includes real-time responsiveness in various operations such as transformation, rotation, scaling, duplication, undoing, as well as physical and geometric editing. To our knowledge, VR-GS represents the first interactive physics-based GS editor.
G3. Unified Framework
VR-GS exemplifies a holistic framework, skillfully integrating both rendering and simulation within a single unified system. Our design philosophy focuses on harmonizing various 3D representation formats, utilizing a blend of thoughtful algorithms and design principles. This integrated approach guarantees a seamless user experience, adhering to the concept of “what you see is what you simulate”(Müller et al., 2016).
G4. User-Centric Design
Essential to a user-friendly VR system is the integration of a flexible model or simulation that supports intuitive and task-oriented user interaction (Stanney et al., 2003). VR-GS is designed to provide this intuitive control, along with responsive feedback, ensuring a natural and seamless interaction experience. The system’s operations and outputs are meticulously aligned with user expectations and commonsense understandings (Hale and Stanney, 2014; Gabbard et al., 1999). Furthermore, VR-GS’s scene in-painting capability, which effectively fills gaps resulting from object removal, is a testament to its commitment to meeting user expectations.
3.2. Design Decisions
Our design decisions align closely with the design goals.
D1. Physics-Based Generative Dynamics
The VR-GS system leverages physics-based simulation techniques to produce realistic dynamics and facilitate 3D shape editing. Unlike methods that reconstruct motion from time-dependent datasets or utilize generative AI to drive Gaussian models, our system adheres to physical principles.
D2. Real-Time Editing and Manipulation
To achieve both real-time performance and physically realistic editing and manipulation, we implement a reduced representation model and parallel-friendly algorithms within our frame budget. While the per-Gaussian-based discretization as done in (Xie et al., 2023) offers detailed dynamics, its time cost is impractical for our system’s needs. Instead, we utilize a tetrahedral mesh embedding reconstructed from Gaussian kernels to drive GS motion. We adopt XPBD with finite-element constraints to achieve real-time simulation paradigm prevalent in gaming and VR applications.
D3. Unified Simulation and Rendering
Our system employs a hybrid approach, combining cage mesh and Gaussian kernels, to achieve real-time performance. To enhance alignment between these two representations, we focus on integrating their functionalities. The deformation gradient field from the cage mesh is adept at embedding a volumeless point cloud. However, as each Gaussian kernel resembles a rotated ellipsoid, it can lead to spiky artifacts during large deformation, a challenge highlighted in (Xie et al., 2023). To mitigate this, we employ a two-level interpolation strategy: initially embedding each Gaussian within an individual tetrahedron, followed by embedding these tetrahedra within the cage mesh to elevate the smoothness of the deformation field.
D4. Natural User Control and Prepossession
Diverging from traditional 2D interface inputs, VR-GS harnesses 3D input mechanisms for interacting with 3D objects. Moreover, the system incorporates object segmentation and scene inpainting functionalities to seamlessly fill gaps created by object removal in the GS training process. This integration ensures a consistent and coherent user experience.
4. Method
4.1. Gaussian Splatting
Gaussian splatting, proposed by Kerbl et al. (2023), is an explicit 3D representation to encapsulate 3D scene information using a set of 3D anisotropic Gaussian kernels, each with learnable mean $\mu$, opacity $\sigma$, covariance matrix $\Sigma$, and spherical harmonic coefficients $\mathcal{C}$. Spherical harmonics (SH) are expansions of the view-dependent color function defined on the unit sphere.
To render a view, the 3D splats are projected to 2D screen space ordered by their $z$-depth. The color $C$ of a pixel is computed by $\alpha$-blending of these 2D Gaussians from near to far:
(1) $C=\sum_{i\in{N}}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$
where $c_{i}$ is the evaluated color by SHs viewed from the camera to the kernel’s mean. $\alpha_{i}$ is the product of the kernel’s opacity and 2D Gaussian weight evaluated at the pixel coordinate. Leveraging a differentiable implementation, the rendering loss towards the ground truth image can be backpropagated to Gaussians’ parameters for optimization.
In contrast to traditional NeRFs based on implicit scene representations, GS provides an explicit representation that can be seamlessly integrated with post-processing manipulations, such as animating and editing. Moreover, the efficient rasterization and superior rendering quality of 3D Gaussians facilitate their integration with VR.
4.2. VR-GS Assets Preparation
Each 3D asset in our VR system consists of a high-fidelity 3D GS reconstruction and an envoloping simulatable tetrahedral mesh in a moderate resolution to enable real-time physics-aware dynamics. These preparations are conducted offline before immersive and interactive editings within our VR environment, which includes:
  • Segmented GS Generation: we support interactions with individual objects in a large scene, achieved by segmented GS reconstructions.
  • Inpainting: The occluded parts between the objects and their supporting planes usually have no texture. We inpaint 3D GS representations leveraging a 2D inpainting technique.
  • Mesh Generation: A simulation-ready tetrahedral mesh is generated for each object.
Segmentation
The segmented GS is constructed during the GS reconstruction. We first generate 2D masks on the multi-view RGB images by utilizing a 2D segmentation model (Cheng et al., 2023). Each segmented part is assigned a different color that is consistent across different views. Subsequently, we enhance the scene representation by integrating three additional learnable RGB attributes into the 3D Gaussian kernels. During the reconstruction process, each 3D Gaussian kernel will automatically learn what object it belongs to utilizing a segmentation loss function $L_{\text{seg}}$:
(2) $L_{\text{seg}}=L_{1}(M_{2d},I),$
where $M_{2d}$ represents colored 2D segmentation results and $I$ denotes the rendering of 3D Gaussian kernels with colors replaced by the extended RGB attributes instead of evaluated from SHs. Thus, the total loss becomes
(3) $L_{\text{total}}=(1-\lambda)L_{1}+\lambda L_{\text{SSIM}}+\lambda_{\text{seg}}L_{\text{seg}},$
where $L_{1}$ and $L_{\text{SSIM}}$ are computed between the normally rendered images and the multi-view ground truths. We use $\lambda=0.2$ and $\lambda_{\text{seg}}=0.1$ in all our experiments.
Inpainting
After 3D GS segmentation, we extract all objects separately from the scene. This process of object removal, however, results in the emergence of holes within the regions that are previously occluded. To alleviate the issue, we utilize 2D inpainting tool LaMa (Suvorov et al., 2022) to guide the 3D inpainting of Gaussian kernels. We freeze the Gaussian kernels located outside of the holes and then use an inpainting loss $L_{\text{inpaint}}=L_{1}(I_{\text{inpainted}},I)$ to optimize a Gaussian kernel patch under the guidance of the 2D inpainted images $I_{\text{inpainted}}$ for the current 3D GS rendering $I$. The result of our inpainting is validated in Section 5.
Mesh Generation
Due to our design choice to use mesh-based simulation, we generate a tetrahedral mesh for each group for the segmented 3D Gaussian kernels. These meshes will not be rendered during the interaction but only serve as the media of dynamics. Hence, we do not require high-fidelity meshes, which, in return, boost the performance of our system. To construct a mesh, we first use internal filling proposed by (Xie et al., 2023) to fill particles into the void internal region, which failed to be reconstructed by GS. Then we treat centers of Gaussians as a point cloud and convert it to a voxel representation represented by the VDB format(Museth, 2013). A water-tight surface mesh is extracted from this VDB using marching cubes (Lorensen and Cline, 1987), which is then served as the boundary of the volume to generate a simulation-ready tetrahedral mesh by TetGen (Si, 2015).
Figure 3. Our two-level embedding effectively resolves spiky artifacts. Each Gaussian kernel is embedded into a local tetrahedron. The vertices of the local tetrahedron are independently embedded into the global mesh.
4.3. Unified Framework for Simulation and Rendering
As derived in Xie et al. (2023), Gaussian kernels can be deformed by the deformation map given by the simulation. We follow the Gaussian kinematics of this work but replace the simulation with eXtended position-based dynamics (XPBD) (Macklin et al., 2016) to achieve real-time interactions. We employ the strain energy constraint as the elastic model and adopt the velocity-based damping model in the XPBD framework. To resolve collisions between meshes in real-time, penalty forces based on point-triangle distance constraint are added on close triangles pairs detected by LBVH (Pantaleoni and Luebke, 2010).
In mesh-based simulation, the deformation map is given by a piece-wise linear map, where the deformation gradient is hence constant within each tetrahedron. Given a tetrahedron with the rest-shape configuration $\{\bm{x}_{0}^{0},\bm{x}_{1}^{0},\bm{x}_{2}^{0},\bm{x}_{3}^{0}\}$ and the current configuration $\{\bm{x}_{0},\bm{x}_{1},\bm{x}_{2},\bm{x}_{3}\}$, the deformation gradient is defined as
(4) $\bm{F}=\begin{bmatrix}\bm{x}_{1}-\bm{x}_{0},\bm{x}_{2}-\bm{x}_{0},\bm{x}_{3}-\bm{x}_{0}\end{bmatrix}\begin{bmatrix}\bm{x}_{1}^{0}-\bm{x}_{0}^{0},\bm{x}_{2}^{0}-\bm{x}_{0}^{0},\bm{x}_{3}^{0}-\bm{x}_{0}^{0}\end{bmatrix}^{-1},$
where the inverse of the rest-shape basis can be computed pre-simulation. The mean and the covariance matrix of the deformed Gaussian kernel inside this tetrahedron is given by
(5) $\bm{\mu}=\sum_{i}w_{i}\bm{x}_{i},\quad\bm{\Sigma}=\bm{F}\bm{\Sigma}_{0}\bm{F}^{T},$
where $w_{i}$ is the barycentric coordinates of the initial center $\bm{\mu}_{0}$ in the rest-shape configuration, and $\bm{\Sigma}_{0}$ is the initial covariance matrix (Xie et al., 2023). The deformed Gaussian kernels can be directly rendered by the point splatting procedure. However, the direct embedding of Gaussian centers cannot guarantee that every ellipse shape is completely enveloped by some tetrahedron inside the simulation mesh. As shown in Figure 3, this can lead to spiky artifacts.
Observe that kernels that are completely inside the tetrahedron will always be inside. This motivates us to propose a two-level embedding procedure:
  • Local embedding: we independently envelope every Gaussian kernel by an as-tight-as-possible tetrahedron. There is no connectivity between these local tetrahedra.
  • Global embedding: we embed the vertices of local tetrahedra into the global simulation mesh.
As the global mesh is deformed by the simulation, the vertices of local tetrahedra are kept inside the boundary, driving the kinetic evolution of Gaussian kernels. A local tetrahedron could overlap with multiple global tetrahedra. The deformation map on it can be understood as the average of the surrounding global tetrahedra, hence eliminating sharp, spiky artifacts, as validated in Section 5.
Shadow Map
While the original global illumination of the scene can be accurately learned and baked by the spherical harmonics on each Gaussian, the shadow will no longer be aligned with the object when it is moving or deforming. The introduction of the shadow map (Wanger et al., 1992) into the GS framework can enhance the immersive experience in VR. More importantly, it can guide human perception of the spatial relationships between objects: during manipulation, users will rely on the shadow to determine the distance between objects. The shadow map is a fast real-time algorithm and is well-aligned with the GS rasterization pipeline. We follow Equation 1 to estimate the depth map from the light source and test the visibility for each Gaussian using this depth map. The influence of shadow maps is ablated in Section 5.
Figure 4. Inpainting Evaluation. GS struggles to reconstruct occluded surfaces. By leveraging LAMA (Suvorov et al., 2022), we produce 2D inpainted results that guide the 3D scene inpainting, enhancing the realism of the 3D representation.
5. Evaluation
In this section, we assess the essential components of our proposed system, including two-level GS embedding and shadow map.
Two-level Embedding
The two-level embedding is a crucial component in our physical dynamics-aware system, integrating the tetrahedra cage mesh with the embedded Gaussian. In the inset figure, we conduct an ablation study to validate the effectiveness of our two-level embedding approach. Under conditions of extreme stretching or twisting, the naive embedding method, which merely embeds the Gaussian kernel within the closest tetrahedron, leads to severe spiky artifacts. Our two-level embedding strategy addresses this by initially embedding each Gaussian into a localized, independent tetrahedron, followed by embedding the vertices of this tetrahedron into the cage mesh. The deformation of each Gaussian kernel is determined by averaging the deformations at these vertices, resulting in a smoother deformation gradient field than the naive approach and significantly reducing spiky artifacts.
Inpainting
With the guidance of 2D segmentation image results, our system achieves object-level 3D scene reconstruction, facilitating convenient post-physics-based manipulation. However, GS is limited to reconstructing surfaces visible in the provided multi-view training images. Consequently, some object and background areas, unseen in these images, remain unreconstructed by GS, leading to ’black hole’ like artifacts when foreground objects are moved. To address this, our system integrates the object segmentation mask with LAMA (Suvorov et al., 2022), a 2D inpainting model, to generate inpainted multi-view images. These images then guide the fine-tuning and inpainting of our 3D GS scene, yielding a more complete and realistic user experience, as evidenced in Figure 4.
Shadow Map
Furthermore, we take advantage of the shadow map to add extra time-dependent shadow into the GS scene, as shown in 5. Original GS models shadows as fixed surface textures, and thus fails to provide dynamic shadows as objects move around the scene. VR-GS allows users to choose the parameters, e.g. position and direction, of the light source to reproduce the lighting setting of the original scene, leading to a more realistic VR environment.
6. Experiment
In this section, we benchmark our VR-GS system’s simulation performance against other leading methods in GS and NeRF manipulation. Additionally, we present immersive and interactive demos within our VR-GS system. Our prototype, developed on Unity with a CUDA-implemented plugin, is tested using a Quest Pro Head-Mounted Display (HMD) and corresponding controllers, on a high-performance Intel(R) Core(TM) i9-14900KF CPU with 32GB memory and an NVIDIA GeForce RTX 4090 GPU.
Figure 5. Shadow Map Evaluation. GS traditionally models shadows as static surface textures. Our approach, employing a shadow map, generates dynamic shadows for a more immersive experience.
6.1. Performance
Example
PAC-
NeRF
Phys-
Gaussian
VR-GS
Stool 0.750 0.112 0.017
Chair 0.813 0.219 0.022
Materials 0.625 0.39 0.021
We compared VR-GS against two state-of-the-art NeRF/GS physics-based manipulation methods: PAC-NeRF (Li et al., 2023a) and PhysGaussian (Xie et al., 2023). PAC-NeRF primarily concentrates on estimating material parameters from multi-view videos in reconstructed NeRF scenes. Although it offers novel dynamic generation capabilities, the resulting visuals often fall short in rendering quality. In contrast, PhysGaussian, leveraging GS, produces superior photorealistic dynamics. However, the Material Point Method (MPM) used by both systems can limit real-time performance in extensive scenes. For a fair comparison, we standardize the frame time ($\Delta t_{\text{frame}}=1/25$ sec) and the simulation step time ($\Delta t_{\text{step}}=0.0001$ sec) across all methods. For VR-GS, we set the XPBD iteration per substep to $1$, as the small $\Delta t_{\text{step}}$ sufficiently resolves the dynamics. As depicted in Figure 6 and the inset table, VR-GS not only matches the visual quality of PhysGaussian but also outperforms PAC-NeRF in clarity and realism. Crucially, our XPBD-based simulation framework allows for significantly faster simulations, making VR-GS ideal for real-time, physics-aware interactions in VR.
Figure 6. Visual Quality Comparison. Our method synthesizes competitive visual results compared to PhysGuassian (Xie et al., 2023) and significantly outperforms PAC-NeRF (Li et al., 2023a).
6.2. VR Demos
Figure 7. Fox, Bear, and Horse. Our proposed VR-GS allows users to manipulate 3D GS at a real-time rate with physically plausible response.
Figure 8. Timing Breakdown of Demos in Section 6.2.
Our VR-GS system’s capability to replicate real-world scenarios is showcased through several representative demos. These demos, employing GS and our dynamic simulation model, illustrate the system’s interactive and immersive qualities. Detailed simulation setups and a timing breakdown for these demos are provided in Table 1 and Figure 8.
Fox, Bear, and Horse Manipulation
VR-GS empowers users to edit 3D Gaussian kernels in a user-friendly manner, utilizing our XPBD-based physics engine for dynamic manipulation. We demonstrate this through three examples: a fox, a bear, and a horse, reconstructed from the Instant-NGP dataset (Müller et al., 2022), the Instruct-NeRF2NeRF (Haque et al., 2023), and the Tanks and Temples dataset (Knapitsch et al., 2017), respectively. Each object’s Gaussian kernels are segmented from their scenes and converted into tetrahedron meshes, allowing for physics-based post-manipulation. The interactions, as depicted in Figure 7, exhibit lively responses to user inputs within the 3D environment.
Ring Toss
In this scenario, we demonstrate the system’s ability to seamlessly integrate new virtual objects into pre-existing scenes. A room scene, reconstructed from our collected real-world data, and a virtual ring, modeled in Blender, are used to create a ring toss game. The dining table and objects on it serve as collision boundaries for the ring. This interaction, shown in Figure 9, highlights VR-GS’s capacity to blend reconstructed and virtual objects in a unified framework, creating an immersive VR experience.
Table 1. Parameters and Timings of Demos in Section 6.2.
Example #Gaussians #Verts. #Iter* CD* FPS
( Figure 7) Fox 304,875 28,205 1 10 161.2
( Figure 7) Bear 2,525,891 60,590 10 10 36.3
( Figure 7) Horse 1,295,223 10,611 5 10 115.4
( Figure 9) Ring Toss 956,116 17,513 20 10 39.0
( Figure 10) Toy Collection 2,555,482 46,810 5 5 19.3
( Figure 4) Box Moving 1,769,412 1,434 10 1 73.8
(Figure 1) Animal Crossing 1,031,191 44,969 15 10 33.5
    #Iter: XPBD iteration count; CD: collision detection per step.
Figure 9. Ring Toss. Utilizing our VR-GS system, participants can engage in an interactive ring toss game simulated within authentic environments.
Toy Collection
To further assess the system’s efficiency and efficacy with multiple interactive objects, we reconstructed a living room scene and a set of plush toys, the latter partly synthesized using LucidDreamer (Liang et al., 2023). This setup, illustrated in Figure 10, allows users to realistically interact with the toys, showcasing VR-GS’s ability to handle complex scenarios involving a mix of 3D GS objects derived from real and synthesized environments.
Animal Crossing
The Animal Crossing background scene is the same as the Ring Toss scene. All four animal models sitting on chairs are reconstructed from Toy Collection. VR-GS permitted free deformation of the animal models, enabling users to freely interact with them. As demonstrated in Figure 1, this setup leveraged our sophisticated generative dynamics, resulting in objects reacting to impacts in a manner that closely mimics real-world physics.
Figure 10. Toy Collection. We place all the toys on the sofa, and then move them into the basket. After shaking the basket, all the toys fly out in a realistic way.
7. User Study
We conducted a user study with 10 participants: 2 professionals (P1–P2) with 5 and 7 years of experience in 3D VFX software (including Houdini and Blender), and 8 novices (P3–P10), with P2 also having 2 years of VR experience. Our study used the hardware specified in the experimental section and included the Fox, Ring Toss, and Toy Collection demos. Participants received a 10-minute video tutorial on our system before completing two tasks. Post-task, they answered a questionnaire covering usability (System Usability Scale (Bangor et al., 2009)) and subjective feedback on individual (M1-M5: object manipulation, inpainted scene, illumination, realistic dynamics, physics-based placement) and overall system features (M6-M9: easiness to use, latency satisfaction, functionality, overall satisfaction), as shown in Section 7.2. Additionally, we conducted a 20-minute semi-structured interview for more in-depth feedback.
7.1. Tasks
Task 1: Goal-directed (30 minutes).
The participants were required to pet a fox (Figure 7) and complete two VR games developed from our system, including ring toss (Figure 9) and toy collection (Figure 10). Each subtask is a 10-minute session. More specifically, in the ring toss task, users are required to throw rings over the objects on the table, and in the toy collection task, users are required to collect messy toys from the living room into a basket. Task 1 has been specifically designed to enable users to interact comprehensively with the primary features of our system. The order of the three subtasks was counterbalanced by a Latin square design to minimize possible learning effects.
Task 2: Open-ended (30 minutes).
The participants are asked to edit the scene and generate dynamics freely with our prototype system, which includes geometry-based editing, physics-based editing, transform, rotation, duplicating, undoing, rescaling, etc.
7.2. Results and Discussion
The inset figure summarizes the subjective feedback about the individual features and system features including easiness to use, latency satisfaction, system functionality, and overall satisfaction. Overall, our system has received positive opinions. Among various features, users spoke highly of the placement of virtual content. Transforming objects without physics rules resulted in floating objects, while in our system, all virtual content placements, such as putting toys onto a sofa, are followed by physics rules and are consistent with the user’s understanding and convention. P7 commented, ”I love this a lot. I have a dog. Petting the fox is really like what I did to my dog at home. I felt so real. Moreover, the placement of the toy is awesome. After they all fell to the basket, the basket even shook for a while” The VR-GS system’s realistic lighting and physics-based dynamics create an immersive experience. As a professional user, P2 spoke highly of high-fidelity generative dynamics and illumination. ”Those motions of virtual content are just like what I see in the physical world. I can’t wait to take photos of my own house and then put them in VR with my video game character.” Users rated the system highly on ease of use (4.6/5) and overall satisfaction (4.8/5). With a System Usability Scale (SUS) score of 83.5, VR-GS is classified as ’excellent’ according to (Bangor et al., 2009).
8. Conclusion and Future Work
This paper has presented VR-GS, a physical dynamics-aware interactive Gaussian Splatting system in VR, particularly first addressing the challenge of editing real-time high-fidelity virtual content. By leveraging the advancements in Gaussian Splatting, VR-GS bridges the quality gap traditionally observed between machine-generated and manually created 3D content. Our system not only enhances the reality and immersion via physically-based dynamics, but also provides fine-grained interaction and manipulation controllablity.
Although all the study participants appreciated the efficiency and effectiveness of our proposed VR-GS, it still remains to be improved. Firstly, rendering high-fidelity Gaussian kernels in VR is computationally demanding. As a result, rendering generative dynamics in a large scene with 2K resolution might lead to potential latency issues in our system. Secondly, we will extend our system to fluid simulation as a future work, a significant aspect of simulation and editing in realistic scenarios. In future iterations, we aim to incorporate a broader range of materials, such as fluid and cloth, to enhance the system’s capabilities.